# GUIDE: Guidance-based Incremental Learning with Diffusion Models

### 06 Mar 2024 v1

@@ 6.2 Alternative variants of guidance
$\begin{split}{\hat{\epsilon}}_{\theta-1}(\mathbf{x}_{t},t,y_{i-1})&=\epsilon_{\theta-1}(\mathbf{x}_{t},t,y_{i-1})\\ &{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}-}s\nabla_{\mathbf{x}_{t}}{\ell}\left(f_{\phi_{i}}(y|\hat{\mathbf{z}}_{0}(\mathbf{x}_{t})),{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}y_{i-1}}\right).\end{split}$ (17)
In both approaches, we increase the diversity of rehearsal samples under frozen or continually trained classifiers. This variation resembles the buffer-based method of Gradient Sample Selection (GSS) (Zaeemzadeh et al., 2019), which seeks to maximize the diversity of the samples stored in the memory buffer.
@@
We evaluate the performance of each proposed variant and present the results of those experiments on CL benchmarks in Tab. 3. We can observe that thanks to the improved quality of rehearsal samples, approaches integrating guidance towards selected classes achieve higher overall performance. In simpler scenarios, increasing the diversity of generated samples can also yield slight improvement, while in more complex settings, its performance is comparable to the baseline. Nevertheless, in all evaluated scenarios, the proposed guidance towards forgotten examples outperforms competing approaches.
@@ 7 Discussion and limitations
In this work, we focus on the continual learning of a classifier given a continually trained diffusion model. In our experiments, we train the diffusion model with simple generative replay – a potentially suboptimal solution. As presented in Appendix B, the significant drawback of our method is the forgetting happening in the diffusion model itself. Our initial results suggest that because of the known issue of sample deficiency from low-density regions (Sehwag et al., 2022) forgetting in diffusion models manifests itself as a degradation of the diversity of generated samples, which has an important effect on the quality of rehearsal samples. We discuss this issue briefly in Appendix D.1. Nevertheless, the forgetting of diffusion models is an open research question.
@@ 8 Conclusion
9 Impact statement
This paper presents work whose goal is to advance the field of continual machine learning. There are many potential societal consequences of our work, many of which are generic to the machine learning field in general, i.e., the DGR algorithm considered in this paper will reflect the biases present in the dataset. Hence, it is important to exercise caution when using this technique in applications where dataset biases could lead to unfair outcomes for minority and/or under-represented groups. In our case, this especially concerns training a diffusion model used for replay, in which simple random sampling can result in a different data distribution than the original dataset.
From the computational point of view and energy consumption, it is essential to find efficient strategies for updating existing diffusion models to new tasks. For instance, the DALL-E 2 model was trained using approximately 650 million images (Ramesh et al., 2022), while the open-source implementation of this method has reported a training time of 56 days using TPU c4 hardware (Dayma et al., 2021). Such computational resources are beyond the reach of many academic institutions, smaller companies, and independent researchers. Our studies on the continual learning of DDPMs contribute to the development of versatile and dynamically evolving generative models with potential applications in a broad spectrum of domains.
Appendix A Impact of classifier scale $s$ hyperparameter on our method
To measure the effect of changing the strength of guidance scale parameter $s$ on the effectiveness of our method, we sweep over each value in $[0.1,0.2,0.5,1.0]$ for each evaluated benchmark. We present the results in Tab. 4. We see that with this hyperparameter properly tuned, we are able to improve the results of our method further. Nonetheless, the results we achieved show that our method works well with different values of this hyperparameter.
Table 4: Effect of changing gradient scale $s$ on GUIDE. By tuning this hyperparameter we can further improve the results.
@@ Appendix A Impact of classifier scale $s$ hyperparameter on our method
Scale $s$ CIFAR-10 CIFAR-100 CIFAR-10 CIFAR-100
$T=2$ $T=5$ $T=5$ $T=10$ $T=2$ $T=5$ $T=5$ $T=10$
0.0 77.43 $\pm$ 0.60 56.61 $\pm$ 1.85 28.25 $\pm$ 0.22 15.90 $\pm$ 1.00 26.32 $\pm$ 0.90 43.79 $\pm$ 0.41 68.70 $\pm$ 0.65 80.38 $\pm$ 1.34
0.1 80.66 $\pm$ 0.44 59.56 $\pm$ 0.52 31.63 $\pm$ 0.81 18.28 $\pm$ 0.93 18.81 $\pm$ 0.48 39.78 $\pm$ 0.92 63.87 $\pm$ 1.35 78.01 $\pm$ 0.59
0.2 81.29 $\pm$ 0.75 61.30 $\pm$ 0.10 37.51 $\pm$ 1.23 22.68 $\pm$ 0.30 14.79 $\pm$ 0.36 34.98 $\pm$ 0.13 54.52 $\pm$ 1.45 70.68 $\pm$ 0.23
0.5 78.30 $\pm$ 0.47 63.12 $\pm$ 0.45 41.66 $\pm$ 0.40 25.48 $\pm$ 1.16 20.03 $\pm$ 1.85 29.25 $\pm$ 1.15 44.30 $\pm$ 1.10 64.06 $\pm$ 1.53
1.0 73.46 $\pm$ 0.81 62.56 $\pm$ 0.52 39.55 $\pm$ 0.19 26.13 $\pm$ 0.29 30.96 $\pm$ 1.93 24.05 $\pm$ 0.72 44.03 $\pm$ 0.34 60.54 $\pm$ 0.82
Moreover, in Fig. 6, we present random rehearsal samples from our CIFAR-10 setup with 2 tasks generated from the same random noise in each column. We observe the effect of classifier scale parameter $s$ on the quality of rehearsal samples in GUIDE. If we set the scale to be too large, we observe significant degradation in the quality of generations. Hence, in Tab. 4, we observe a drop in performance on setups where the scale is too large.
Figure 6: Random rehearsal samples of GUIDE generated by class-conditional diffusion model trained on the first task of CIFAR-10/2 setup. If we set the gradient scale parameter $s$ too large, we observe a significant drop in the quality of samples.
@@ Appendix B Continual learning with ground-truth diffusion
In all of our setups presented in the main part of the article, we trained the diffusion model alongside the continually trained classifier. In this experiment, we study the effect of GUIDE under the assumption of a ground-truth sampling method from the previous tasks. We independently train five diffusion models on all of the data presented up to the $i^{th}$ tasks in the CIFAR100/5 tasks scenario. With those models, we compare our method with the baseline sampling method and continual upper-bound where the classifier has full access to all training data samples from current and previous tasks. We depict the results of this comparison in Fig. 7, showing that the random sampling of rehearsal samples, even from the ground-truth diffusion model, leads to a significant drop in performance. At the same time, our method significantly reduces the difference with the soft-upper bound defined by the continual-joint training with infinite buffer size. Interestingly, our fully continual learning setup still outperforms the baseline sampling approach rehearsed with a ground-truth diffusion.
@@
Figure 7: Average accuracy on all classes seen so far after $i^{th}$ task. We highlight with asterisks ()($\ast$)( ∗ ) methods trained with ground-truth diffusion model. We can observe that under ground-truth diffusion model GUIDE approaches the soft upper-bound defined by the continual join training.
Appendix C Training details and hyperparameters
C.1 Architectures
@@
In our experiments, we follow the definitions of class-conditional diffusion model architectures described in Dhariwal & Nichol (2021). We use the same diffusion model architecture for each experimental setup. As a classifier, we use the same ResNet18 architecture as the GDumb method (Prabhu et al., 2020) in each setup, with preactivation enabled, meaning that we place the norms and activations before the convolutional or linear layers.
C.2 Diffusion models training
For setups with 5 tasks, we train models for 50000 steps per task, whereas for the CIFAR-10 with 2 tasks and CIFAR-100 with 10 tasks, we train for 100000 steps per task. Across all setups, we maintain consistent hyperparameters: the AdamW optimizer is used with a batch size of 256 and a learning rate of 0.0002. Our diffusion models are trained using linear noise schedulers over 1000 steps, featuring U-Net-based architectures with 3 residual blocks.
C.3 Classifiers training
We train models with an SGD optimizer. During the first task, we set the learning rate to 0.1, and in each subsequent task to 0.01. We set the batch size to 256.
In each iteration of training, we split the batch equally between all tasks encountered so far. We generate the same number of rehearsal samples from each encountered task, then we shuffle them along with real data samples from the current task and apply the augmentations on both generated and real data samples.
C.4 Augmentations
@@
For all datasets, we define the same set of image augmentations, which include cropping, rotating, flipping, and erasing. We also apply a transformation to the brightness, contrast, saturation, and hue, followed by the normalization to the [-1, 1] range. During training, we apply the same augmentations for both rehearsal samples and data from the current task.
Appendix D Analysis of forgetting in diffusion models
D.1 Coverage of data manifold

### 31 May 2024 v2

@@ 6.2 Alternative variants of guidance
${\hat{\epsilon}}_{\theta_{i-1}}(\mathbf{x}_{t},t,y_{i-1})=\epsilon_{\theta_{i-1}}(\mathbf{x}_{t},t,y_{i-1}){\color[rgb]{0,0,1}-}s\nabla_{\mathbf{x}_{t}}{\ell}\left(f_{\phi_{i}}(y|\hat{\mathbf{z}}_{0}(\mathbf{x}_{t})),{\color[rgb]{0,0,1}y_{i-1}}\right).$ (14)
Figure 5: Classifier scale impact on forgetting and ability to acquire new information. Up to $s=0.2$, when we increase scale, we reduce the forgetting but also observe a drop in the accuracy on the second task. When we use too large scale $s$, the quality of samples drops significantly, along with the accuracy on the previous task. We further present this effect in Appendix A.
In both approaches, we increase the diversity of rehearsal samples under frozen or continually trained classifiers. This variation resembles the buffer-based method of Gradient Sample Selection (GSS) [53], which seeks to maximize the diversity of the samples stored in the memory buffer.
@@
We evaluate the performance of each proposed variant and present the mean and standard deviation of those experiments on CL benchmarks in Tab. 3, which are calculated for 3 random seeds. We can observe that thanks to the improved quality of rehearsal samples, approaches integrating guidance towards selected classes achieve higher overall performance. In simpler scenarios, increasing the diversity of generated samples can also yield slight improvement, while in more complex settings, its performance is comparable to the baseline. Nevertheless, in all evaluated scenarios, our proposed guidance towards forgotten examples outperforms competing approaches.
@@ 7 Discussion and limitations
In this work, we focus on the continual learning of a classifier given a continually trained diffusion model. In our experiments, we train the diffusion model with simple generative replay – a potentially suboptimal solution. As presented in Appendix B, the significant drawback of our method is the forgetting happening in the diffusion model itself. Our initial results suggest that because of the known issue of sample deficiency from low-density regions [40] forgetting in diffusion models manifests itself as a degradation of the diversity of generated samples, which has an important effect on the quality of rehearsal samples. We discuss this issue briefly in Appendix E. Nevertheless, the forgetting of diffusion models is an open research question. An important drawback of all generative replay approaches is the computational burden associated with training and sampling from the diffusion model. This is also true for our method, which is computationally expensive. However, GUIDE is only slightly more computationally expensive than standard DGR with diffusion and significantly less expensive than a recent state-of-the-art solution – DDGR, while yielding better results. Moreover, we evaluate additional speed-up techniques that mitigate this burden. Detailed analysis is provided in Appendix G.
@@ 8 Conclusion
Appendix A Impact of classifier scale hyperparameter on our method
To measure the effect of changing the strength of guidance scale parameter $s$ on the effectiveness of our method, we sweep over each value in $[0.1,0.2,0.5,1.0]$ for each evaluated benchmark. We present the mean and standard deviation of results in Tab. 4 calculated for 3 different values of random seed. We see that with this hyperparameter properly tuned, we are able to improve the results of our method further. Nonetheless, the results we achieved show that our method works well with different values of this hyperparameter.
Table 4: Effect of gradient scale $s$ on GUIDE. By tuning this hyperparameter, we can improve the results further.
@@ Appendix A Impact of classifier scale hyperparameter on our method
Scale $s$ CIFAR-10 CIFAR-100 ImageNet100-64 CIFAR-10 CIFAR-100 ImageNet100-64
$T=2$ $T=5$ $T=5$ $T=10$ $T=5$ $T=2$ $T=5$ $T=5$ $T=10$ $T=5$
0.0 77.43 $\pm$ 0.60 56.61 $\pm$ 1.85 28.25 $\pm$ 0.22 15.90 $\pm$ 1.00 23.92 $\pm$ 0.92 26.32 $\pm$ 0.90 43.79 $\pm$ 0.41 68.70 $\pm$ 0.65 80.38 $\pm$ 1.34 54.44 $\pm$ 0.14
0.1 80.66 $\pm$ 0.44 59.56 $\pm$ 0.52 31.63 $\pm$ 0.81 18.28 $\pm$ 0.93 26.48 $\pm$ 3.79 18.81 $\pm$ 0.48 39.78 $\pm$ 0.92 63.87 $\pm$ 1.35 78.01 $\pm$ 0.59 50.09 $\pm$ 5.04
0.2 81.29 $\pm$ 0.75 61.30 $\pm$ 0.10 37.51 $\pm$ 1.23 22.68 $\pm$ 0.30 31.09 $\pm$ 4.17 14.79 $\pm$ 0.36 34.98 $\pm$ 0.13 54.52 $\pm$ 1.45 70.68 $\pm$ 0.23 44.40 $\pm$ 5.02
0.5 78.30 $\pm$ 0.47 64.47 $\pm$ 0.45 41.66 $\pm$ 0.40 25.48 $\pm$ 1.16 35.82 $\pm$ 0.56 20.03 $\pm$ 1.85 29.25 $\pm$ 1.15 44.30 $\pm$ 1.10 64.06 $\pm$ 1.53 35.80 $\pm$ 0.32
1.0 73.46 $\pm$ 0.81 62.56 $\pm$ 0.52 39.55 $\pm$ 0.19 26.13 $\pm$ 0.29 39.07 $\pm$ 1.37 30.96 $\pm$ 1.93 24.05 $\pm$ 0.72 44.03 $\pm$ 0.34 60.54 $\pm$ 0.82 27.60 $\pm$ 3.29
Moreover, in Fig. 6, we present random rehearsal samples from our CIFAR-10 setup with 2 tasks generated from the same initial noise in each column. We observe the effect of classifier scale parameter $s$ on the quality of rehearsal samples in GUIDE. If we set the scale to be too large, we observe significant degradation in the quality of generations. Hence, in Tab. 4, we observe a drop in performance on setups where the scale is too large.
Figure 6: Sample rehearsal examples of GUIDE generated by class-conditional diffusion model trained on the first task of CIFAR-10/2 setup. If we set the gradient scale parameter $s$ too large, we observe a significant drop in the quality of samples.
@@ Appendix B Continual learning with ground-truth diffusion
In all of our setups presented in the main part of the article, we trained the diffusion model alongside the continually trained classifier. In this experiment, we study the effect of the rehearsal sampling method proposed in GUIDE under the assumption of a ground-truth sampling method from the previous tasks. We independently train four diffusion models on all of the data presented up to the $i^{th}$ task in the CIFAR100/5 tasks scenario. With those models, we compare our method with the baseline sampling method and continual upper-bound where the classifier has full access to all training data samples from current and previous tasks. We depict the results of this comparison in Fig. 7, calculated for 3 random seeds. We report mean and standard deviation as error bars. Results indicate that the random sampling of rehearsal samples, even from the ground-truth diffusion model, leads to a significant drop in performance. At the same time, our method significantly reduces the difference with the soft-upper bound defined by the continual-joint training with infinite buffer size. Interestingly, our fully continual learning setup still outperforms the baseline sampling approach rehearsed with a ground-truth diffusion.
@@
Figure 7: Average accuracy on all classes seen so far after $i^{th}$ task. We highlight with asterisks ()($\ast$)( ∗ ) methods trained with ground-truth diffusion model. We can observe that under ground-truth diffusion model GUIDE approaches the soft upper-bound defined by the continual joint training.
Appendix C Pseudocode of training procedure in GUIDE
To clarify and enhance understanding of our continual training process, this section includes pseudocode for the primary components involved in our method. Alg. 1 presents a sampling of a single rehearsal example during the continual training of a classifier in GUIDE. In Alg. 2, we present a complete end-to-end continual training of both diffusion model and classifier in GUIDE.
Since the decision boundary changes during the training of a classifier, we sample new rehearsal examples according to our sampling method progressively once every $N_{g}$ batches. It ensures that replayed examples are located close to the decision boundary of a classifier at the given moment. Simultaneously, hyperparameter $N_{g}$ allows us to significantly speed up our method without a noticeable drop in accuracy, which we describe further in Appendix G. Since we do not use classifier guidance during the generation of a dataset for diffusion training, we sample rehearsal examples all at once and merge them with real data samples.
In the pseudocode, consistent with the terminology used in the main text, the current classifier is denoted as $f_{\phi_{i}}(y|\mathbf{x})$ and the previous diffusion model as $\epsilon_{\theta_{i-1}}(\mathbf{x}_{t},t,y)$. All hyperparameters, such as the number of training steps or the number of denoising steps, are listed in Appendix D.
Algorithm 1 Rehearsal sampling in GUIDE during training on task $i$
$\mathcal{C}_{i-1}$: classes from all previous tasks, $\mathcal{C}_{i}$: classes from current task, $s$: gradient scale, $T$: number of denoising steps
$t\leftarrow T$
$y_{i-1}\sim\mathcal{U}(\mathcal{C}_{i-1})$
$\mathbf{x}_{t}\sim\mathcal{N}(0,\mathbf{I})$
while $t>0$ do
     $\hat{\mathbf{z}}_{0}(\mathbf{x}_{t})\leftarrow\frac{\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta_{i-1}}(\mathbf{x}_{t},t,y_{i-1})}{\sqrt{\bar{\alpha}_{t}}}$
     $y_{i}\leftarrow\underset{c\in\mathcal{C}_{i}}{\mathrm{argmax}}f_{\phi_{i}}(y=c|\hat{\mathbf{z}}_{0}(\mathbf{x}_{t}))$
     ${\hat{\epsilon}}_{\theta_{i-1}}(\mathbf{x}_{t},t,y_{i-1})\leftarrow\epsilon_{\theta_{i-1}}(\mathbf{x}_{t},t,y_{i-1})+s\nabla_{\mathbf{x}_{t}}{\ell}\left(f_{\phi_{i}}(y|\hat{\mathbf{z}}_{0}(\mathbf{x}_{t}),y_{i}\right)$
     $\mathbf{x}_{t-1}\leftarrow\sqrt{\bar{\alpha}_{t-1}}\hat{\mathbf{z}}_{0}(\mathbf{x}_{t})+\sqrt{1-\bar{\alpha}_{t-1}}{\hat{\epsilon}}_{\theta_{i-1}}(\mathbf{x}_{t},t,y_{i-1})$
     $t\leftarrow t-1$
end while
$\mathbf{x}_{0}$triangleright Generated rehearsal sample from class $y_{i-1}$
Algorithm 2 Continual training in GUIDE
$N_{c}$: number of classifier training steps, $N_{d}$: number of diffusion training steps, $N_{g}$: rehearsal generation interval, $T$: number of tasks, $B$: batch size, $f_{\phi}$: classifier model, $\epsilon_{\theta}$: diffusion model
for $i\in[1,\dots,T]$ do
     $\mathcal{D}_{i}\leftarrow$ real dataset for task $i$
     $\epsilon_{\theta_{i}}\leftarrow\epsilon_{\theta_{i-1}}$
     $f_{\phi_{i}}\leftarrow f_{\phi_{i-1}}$
     $\hat{\mathcal{B}}\leftarrow\emptyset$
     for $n\in[1,\dots,N_{c}]$ do $\triangleright$ Classifier training
         if $i==1$ then
              $\mathcal{B}\leftarrow B$ real samples from $\mathcal{D}_{i}$
         else
              $\mathcal{B}\leftarrow\nicefrac{{B}}{{i}}$ real samples from $\mathcal{D}_{i}$
              if $n\bmod N_{g}==0$ then
                  ^$\hat{\mathcal{B}}\leftarrow\texttt{GUIDE\_sample}(\texttt{num\_samples}=\nicefrac{{B}}{{i}}\times(i-1),f_{\phi_{i}},\epsilon_{\theta_{i-1}})$
              end if
         end if
         $\mathcal{B}_{c}\leftarrow\mathcal{B}\cup\hat{\mathcal{B}}$
         update $f_{\phi_{i}}$ with $\mathcal{B}_{c}$
     end for
     $\hat{\mathcal{D}_{i}}\leftarrow\emptyset$triangleright Construct diffusion dataset
     if $i>1$ then
         $\hat{\mathcal{D}_{i}}\leftarrow\texttt{sample\_without\_guidance}(\texttt{num\_samples}=|\mathcal{D}_{1,\dots,i-1}|,\epsilon_{\theta_{i-1}})$
     end if
     $\mathcal{D}_{d}\leftarrow\mathcal{D}_{i}\cup\hat{\mathcal{D}_{i}}$
     train $\epsilon_{\theta_{i}}$ on $\mathcal{D}_{d}$ for $N_{d}$ steps
end for
Appendix D Training details and hyperparameters
D.1 Diffusion models
In our experiments, we follow the definitions of class-conditional diffusion model architectures described in [8]. Across all setups, we train diffusion models using AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The only hyperparameter that varies between tasks is the number of iterations of training. In all setups, we train models for 100K iterations on the initial task, and then on each subsequent task, we train either for 50K or 100K iterations, depending on the setup. We use DDIM with 250 steps for the generation of rehearsal samples on the ImageNet100 dataset and 1000 denoising steps for every other dataset. The only augmentation that we use during the training of diffusion models is the random horizontal flip.
We present the most important hyperparameters for diffusion models in Tab. 5.
Table 5: Hyperparameters for the training of diffusion models.
CIFAR-10/2 CIFAR-10/5 CIFAR-100/5 CIFAR-100/10 ImageNet100-64/5
Diffusion steps 1000 1000 1000 1000 1000
Noise schedule linear linear linear linear linear
Channels 128 128 128 128 192
Depth 3 3 3 3 3
Channels multiple 1, 2, 2, 2 1, 2, 2, 2 1, 2, 2, 2 1, 2, 2, 2 1, 2, 3, 4
Heads 4 4 4 4
Heads channels 64
Attention resolution 16,8 16,8 16,8 16,8 32,16,8
BigGAN up/downsample
Dropout 0.1 0.1 0.1 0.1 0.1
Batch size 256 256 256 256 100
Learning rate 2e-4 2e-4 2e-4 2e-4 1e-4
Iterations 1st task 100K 100K 100K 100K 100K
Iterations other tasks - 50K 50K 100K 50K
Self-rehearsal denoising steps - 1000 1000 1000 DDIM250
D.2 Classifiers
@@
As a classifier model, we use the same ResNet18 architecture as the GDumb method [34] in each setup, with preactivation enabled, meaning that we place the norms and activations before the convolutional or linear layers. We train models with an SGD optimizer. We list the most important hyperparameters for classifiers in Tab. 6.
@@
For CIFAR-10 and CIFAR-100 datasets, we define the same set of image augmentations, which include cropping, rotating, flipping, and erasing. We also apply a transformation to the brightness, contrast, saturation, and hue, followed by the normalization to the [-1, 1] range. During the training, we apply the same augmentations for both rehearsal samples and real samples from the current task.
For the ImageNet100, we do not apply any data augmentations, but we also normalize the images to the range of [-1, 1].
Table 6: Hyperparameters for the training of classifier models.
CIFAR-10/2 CIFAR-10/5 CIFAR-100/5 CIFAR-100/10 ImageNet100-64/5
Batch size 256 256 256 256 100
Learning rate 1st task 0.1 0.1 0.1 0.1 0.1
Learning rate other tasks 0.01 0.01 0.05 0.05 0.001
Iterations 1st task 5K 5K 10K 10K 20K
Iterations other tasks 2K 2K 2K 2K 20K
Rehearsal denoising steps DDIM50 DDIM50 DDIM100 DDIM100 DDIM50
Rehearsal generation interval 1 5 10 10 15
D.3 Experiments compute resources
In this section, we list the hardware we used for our experiments and training times on each benchmark, both for training of diffusion models in Tab. 7 and classifiers in Tab. 8. Although we use NVIDIA A100 GPUs in our experiments due to the efficiency of training, experiments can also be reproduced on GPUs with smaller memory, as noted in Sec. F.
The total compute used in a project can be estimated from the times presented in tables, considering that each experiment is calculated with three different random seeds for statistical significance. Furthermore, we need to take into account compute used in the development stage of the project.
Table 7: Training times of diffusion models for each benchmark.
Benchmark Time[GPU-hours] GPU used
CIFAR-10/2 5.55 4 x NVIDIA A100 40GB
CIFAR-10/5 20 4 x NVIDIA A100 40GB
CIFAR-100/5 20 4 x NVIDIA A100 40GB
CIFAR-100/10 63.75 4 x NVIDIA A100 40GB
ImageNet100-64/5 88.43 4 x NVIDIA A100 40GB
Table 8: Training times of classifiers for each benchmark.
Benchmark Time[GPU-hours] GPU used
CIFAR-10/2 4.8 1 x NVIDIA A100 40GB
CIFAR-10/5 5.8 1 x NVIDIA A100 40GB
CIFAR-100/5 5.35 1 x NVIDIA A100 40GB
CIFAR-100/10 13.53 1 x NVIDIA A100 40GB
ImageNet100-64/5 30.73 4 x NVIDIA A100 40GB
Appendix E Forgetting in diffusion models
E.1 Coverage of data manifold